1. show all work and formulas
below is a hypothetical data for a sample of british children and their IQ scores
female 9,11,10,13,8,9,11,9
A)what is the average IQ score? show all work
B) what is the sandard deviation for IQ scores?
C)transform the raw scores 11,8 and 9 into Z scores using the sample mean score and standard deviation caculated in previous steps?
D)what percentile for the Raw Score 8 and 9 in the women condition?
E)what is the aw score for a woman whohas a Z score of 1.32?
Answer #1

A) Average score = sum of data values / number of data values

= 80/8

=10

B) Standard deviation = sqrt(summation of(x-xbar)^2/n)

= Sqrt(2.57)

= 1.6

C) z score = (raw score-mean) / standard deviation

z score corresponding to

11 = (11-10)/1.6 = 0.625

8 = (8-10)/1.6 = -1.25

9 = (9-10)/1.6 = -0.625

D) Raw score 8 and 9 shows IQ is below average IQ level.

E) Raw score = (z value*standard deviation) + mean

= (1.32*1.6) + 10

= 12.112